# Foveation Improves Payload Capacity in Steganography

### v1

@@ 1. Introduction
Steganography concerns about hiding data in another medium (Wang et al., 2023). More specifically, our work studies embedding information in images. Steganography is therefore useful in conveying multi-modal information such as labels, scene descriptions, or copyright marks. Growing number of AI-generated content and introduction of AR/VR systems increases its importance further by broadening the application scope (Rezaei et al., 2024).
Figure 1. Visual example of stego and payload recovery using our proposed foveated steganography and RoSteALS (Source: MetFaces (Karras et al., 2020)).:
Our work leverages latent representations (Yılmaz et al., 2024) and a foveated rendering loss (Walton et al., 2022) to increase payload capacity in steganography. With only 2000 training images, we achieve bit accuracy of 99.99% for 40K test set bits. Concerning state of the art in latent methods (Bui et al., 2023), our approach increases the payload capacity from 100 bits up to 500 bits with up to 100% recovery under non-distortion condition, 27.56 db PSNR and 0.26 LPIPS. Our final contribution is the introduction of Metameric Foveated Rendering loss in steganography, which noticeably improves all visual metrics and quality with respect to classic L2 loss.
@@ 2. Method
Considering a message payload $P\in\{0,1\}^{k}$ consisting of $k$ bits and a input image (cover) $I\in\mathbb{R}^{h\times w\times c}$, find two functions $H$ and $R$, to produce an output image (stego), $H(I,P)=I^{\prime}\in\mathbb{R}^{h\times w\times c}$, and output payload, $R(I^{\prime})=P^{\prime}\in\{0,1\}^{k}$. The aim is to reduce the distortion between $I$ and $I^{\prime}$ while maximizing the accuracy between $P$ and $P^{\prime}$.
Figure 2. Our proposed foveated steganography approach (Source: MetFaces (Karras et al., 2020)).:
@@
Our framework approaches this problem as depicted in Figure 2. In the hiding stage, a frozen image encoder, $E$, transforms input image into a latent representation, $E(I)=Z_{i}$. Payload embedder, $F$, creates also a learned representation, $P(I)=Z_{p}$. Together, they are manipulated by the merger, $M$, producing a merged latent, $M(Z_{i},Z_{p})=Z_{m}$, which a frozen image generator, $G$, uses to reconstruct the output image, $G(Z_{m})=I^{\prime}$. Finally, a payload retriever, $R$, extracts the output payload, $R(I^{\prime})=P^{\prime}$. The loss function is defined as the combination of payload and image quality losses, being BCE and Metameric Foveated Rendering (Walton et al., 2022) (defaulted to center) losses respectively. Formally, $\mathcal{L}_{total}=\mathcal{L}_{payload}+\lambda_{i}\cdot\mathcal{L}_{image}=\operatorname{BCE}(P,P^{\prime})+\lambda_{i}\cdot(\operatorname{MetamericLoss}(I,I^{\prime})),$ where $\lambda_{i}$ controls the trade-off between the two losses.
The dataset is a balanced mixture of 2000 training, 400 validation, and 400 test images from MetFaces (Karras et al., 2020) and CLIC datasets (Toderici et al., 2020). For preprocessing, images are randomly cropped and padded to the size of input and normalized as autoencoder requires. Notably, this dataset is much smaller than typical datasets used for the same purpose, but is found sufficient to learn performing 100-bits steganography, within controlled computing resources, about two hours on a single RTX 4090 GPU.
@@ 3. Results and Discussion
The frozen pair of image encoder and image generator to create a high-quality latent representation, is the F4-with-attention version autoencoder from LDM VQGAN series (Rombach et al., 2022). After evaluating empirically, we found its high reconstruction quality is suitable for the embedding process. Compared to other backbones, this one converges slower at payload embedding, but achieves better image quality in the end. Keeping payload embedder as a fully connected layers is sufficient to encode the information after experimentation. For merger, the best performing architecture is adding two convolutional layer sandwiching the sum of image and payload latent, to soften the transition. Finally, ResNet50 was used as payload retriever as a popular and well-studied architecture.
@@
The main results are shown in Table 1. Baselines achieved expanding payload capacities, at various resolutions. At minimal setting, baseline has a bit accuracy of 99.99%, failing to decode only 4 out of 40K test bits. Noticeably, we achieve 100% recovery in the native resolution of benchmark, RoSteALS (Bui et al., 2023), while other settings also all exceed 99.95% compared to benchmark failing to reach 99.5%. Nevertheless, RoSteALS has better perceptual image quality. This is reasonable since it is an augmented version of the baseline which uses larger datasets, incorporates LPIPS in loss, and applies finer-grained optimization in the training.

### v2

@@ 1. Introduction
Steganography concerns about hiding data in another medium (wang2023data). More specifically, our work studies embedding information in images. Steganography is therefore useful in conveying multi-modal information such as labels, scene descriptions, or copyright marks. Growing number of AI-generated content and introduction of AR/VR systems increases its importance further by broadening the application scope (rezaei2024lawa).
Figure 1. Visual example of stego and payload recovery using our proposed foveated steganography and RoSteALS (Source: MetFaces (karras2020training)).:
Our work leverages latent representations (yilmaz2024learned) and a foveated rendering loss (walton2022metameric) to increase payload capacity in steganography. With only 2000 training images, we achieve bit accuracy of 99.99% for 40K test set bits. Concerning state of the art in latent methods (bui2023rosteals), our approach increases the payload capacity from 100 bits up to 500 bits with up to 100% recovery under non-distortion condition, 27.56 db PSNR and 0.26 LPIPS. Our final contribution is the introduction of Metameric Foveated Rendering loss in steganography, which noticeably improves all visual metrics and quality with respect to classic L2 loss.
@@ 2. Method
Considering a message payload $P\in\{0,1\}^{k}$ consisting of $k$ bits and a input image (cover) $I\in\mathbb{R}^{h\times w\times c}$, find two functions $H$ and $R$, to produce an output image (stego), $H(I,P)=I^{\prime}\in\mathbb{R}^{h\times w\times c}$, and output payload, $R(I^{\prime})=P^{\prime}\in\{0,1\}^{k}$. The aim is to reduce the distortion between II and $I^{\prime}$ while maximizing the accuracy between $P$ and $P^{\prime}$.
Figure 2. Our proposed foveated steganography approach (Source: MetFaces (karras2020training)).:
@@
Our framework approaches this problem as depicted in Figure 2. In the hiding stage, a frozen image encoder, $E$, transforms input image into a latent representation, $E(I)=Z_{i}$. Payload embedder, $F$, creates also a learned representation, $P(I)=Z_{p}$. Together, they are manipulated by the merger, $M$, producing a merged latent, $M(Z_{i},Z_{p})=Z_{m}$, which a frozen image generator, $G$, uses to reconstruct the output image, $G(Z_{m})=I^{\prime}$. Finally, a payload retriever, $R$, extracts the output payload, $R(I^{\prime})=P^{\prime}$. The loss function is defined as the combination of payload and image quality losses, being BCE and Metameric Foveated Rendering (walton2022metameric) (defaulted to center) losses respectively. Formally, $\mathcal{L}_{total}=\mathcal{L}_{payload}+\lambda_{i}\cdot\mathcal{L}_{image}=\operatorname{BCE}(P,P^{\prime})+\lambda_{i}\cdot(\operatorname{MetamericLoss}(I,I^{\prime})),$ where $\lambda_{i}$ controls the trade-off between the two losses.
The dataset is a balanced mixture of 2000 training, 400 validation, and 400 test images from MetFaces (karras2020training) and CLIC datasets (toderici2020workshop). For preprocessing, images are randomly cropped and padded to the size of input and normalized as autoencoder requires. Notably, this dataset is much smaller than typical datasets used for the same purpose, but is found sufficient to learn performing 100-bits steganography, within controlled computing resources, about two hours on a single RTX 4090 GPU.
@@ 3. Results and Discussion
The frozen pair of image encoder and image generator to create a high-quality latent representation, is the F4-with-attention version autoencoder from LDM VQGAN series (rombach2022high). After evaluating empirically, we found its high reconstruction quality is suitable for the embedding process. Compared to other backbones, this one converges slower at payload embedding, but achieves better image quality in the end. Keeping payload embedder as a fully connected layers is sufficient to encode the information after experimentation. For merger, the best performing architecture is adding two convolutional layer sandwiching the sum of image and payload latent, to soften the transition. Finally, ResNet50 was used as payload retriever as a popular and well-studied architecture.
@@
The main results are shown in Table 1. Baselines achieved expanding payload capacities, at various resolutions. At minimal setting, baseline has a bit accuracy of 99.99%, failing to decode only 4 out of 40K test bits. Noticeably, we achieve 100% recovery in the native resolution of benchmark, RoSteALS (bui2023rosteals), while other settings also all exceed 99.95% compared to benchmark failing to reach 99.5%. Nevertheless, RoSteALS has better perceptual image quality. This is reasonable since it is an augmented version of the baseline which uses larger datasets, incorporates LPIPS in loss, and applies finer-grained optimization in the training.